myininaya (myininaya):

integrate (sqrt((x-1)/x^5)) using substitution (no trig substitution)

myininaya (myininaya):

\[\int\limits_{}^{}\sqrt{\frac{x-1}{x^5}} dx\]

OpenStudy (anonymous):

this one is really..........not that tough

OpenStudy (anonymous):

take out sqrt (1/x^4)

OpenStudy (anonymous):

by "take out" i mean we can rewrite the problem as: \[\int\limits_{}^{}\sqrt{1/x^4}* \sqrt{x-1/x}\]

OpenStudy (anonymous):

clever

OpenStudy (anonymous):

typo though right?

OpenStudy (anonymous):

\[\int\frac{1}{x^2}\sqrt{1-\frac{1}{x}}dx\]

OpenStudy (anonymous):

i posted one for you myininaya just in case you have to teach cacl 1 instead of calc 2

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{x^2} \sqrt{\frac{x-1}{x}} dx\] let u=x-1 du=dx if u=x-1=> u+1=x so we have \[\int\limits_{}^{}\frac{1}{(u+1)^2} \cdot \sqrt{\frac{u}{u+1}} du\] let s=u/(u+1) then ds=1/(u+1)^2 du so we have \[\int\limits_{}^{}\sqrt{s} ds\] gj lagrange

OpenStudy (anonymous):

what is wrong with \[u=1-\frac{1}{x}\] \[du=\frac{1}{x^2}dx\] get \[\int\sqrt{u}du\]?

OpenStudy (anonymous):

thats the way i would have done it :)

myininaya (myininaya):

nothing wrong with it

myininaya (myininaya):

lol i do my own thing

myininaya (myininaya):

leave me alone

OpenStudy (anonymous):

more than one way to skin a muskrat. i am waiting for you to prove the product rule from the rule for squares

myininaya (myininaya):

ok zarkon will probably answer before me but i will take a stab at it